OpenStudy (anonymous):

what is the inverse function of the function Y=(2X)/(X+5)

Parth (parthkohli):

In an inverse function, you always replace x with y and y with x, right?

OpenStudy (anonymous):

so would you change the equations to X=(2Y)/X+5)?

OpenStudy (anonymous):

I think that the answer is y= (7/2)x I am not sure. Try using a graphing calculator and then reflect it on the y=x line. You can use online graphing. Good luck. Hope I helped.

OpenStudy (anonymous):

how would you get (7/2)x if you replace the x with y and y with x? @boblovesmath @parthkohli

OpenStudy (anonymous):

I told you I am not very sure of the answer. But what I do know is that you have to cross multiply, then divide by y, and then get common denominators and then add. I ended up with (2+5)/2y = 1/x. As I told you before, I am not very sure. Good luck.

OpenStudy (anonymous):

okay thanks! i'll try that